Find (a) the mean, (b) the median, (c) the mode, and (d) the midrange.



Observers in the Frogwatch Program. The number of observers in the Frogwatch USA


program (a wildlife conservation program dedicated to helping conserve frogs and toads) for the


top 10 states with the most observers is 484, 483, 422, 396, 378, 352, 338, 331, 318, and 302. The


top 10 states with the most active watchers list these numbers of visits: 634, 464, 406, 267, 219,


194, 191, 150, 130, and 114. Compare the measures of central tendency for these two groups of


data

1
Expert's answer
2022-06-06T05:18:23-0400

Sorted Data Set: 


"302, 318, 331, 338, 352, 378, 396, 422, 483, 484"



Sorted Data Set: 


"114, 130, 150, 191, 194, 219, 267, 406, 464, 634"


(a)


"mean_1=\dfrac{1}{10}(302+318+331+338+352"




"+378+396+422+483+484)=380.4"


"mean_2=\dfrac{1}{10}(114+130+150+191+194"




"+219+ 267+406+464+634)=276.9"


"mean_1>mean_2"

(b)


"median_1=\dfrac{352+378}{2}=365"




"median_2=\dfrac{194+219}{2}=206.5"

"median_1>median_2"

(c)


"mode_1=undefined,"

All values appeared just once.



"mode_2=undefined,"

All values appeared just once.



(d)


"midrange_1=\dfrac{302+484}{2}=393"




"midrange_2=\dfrac{114+634}{2}=374"

"midrange_1>midrange_2"